OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@pooja195 @ParthKohli

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@ganeshie8

OpenStudy (er.mohd.amir):

A

OpenStudy (anonymous):

u r sure

OpenStudy (er.mohd.amir):

since 0*0=0 1*1=1 0*1=0 1*0=0

OpenStudy (anonymous):

Choose all answers that are correct.

OpenStudy (anonymous):

and c

OpenStudy (er.mohd.amir):

addition not possible since 1+1=2 not in set

OpenStudy (anonymous):

d

OpenStudy (anonymous):

99999999999.9999999999999999999999% sure

OpenStudy (er.mohd.amir):

subtraction not possible 0-1=-1 not in set

OpenStudy (anonymous):

b

OpenStudy (er.mohd.amir):

division not possible 1/0 and 0/0 not difine

OpenStudy (anonymous):

then

OpenStudy (anonymous):

Choose all answers that are correct.

OpenStudy (anonymous):

a and .......

OpenStudy (er.mohd.amir):

A

OpenStudy (anonymous):

and ....

OpenStudy (anonymous):

u r sure

OpenStudy (anonymous):

can u help me with more

OpenStudy (er.mohd.amir):

i give U reason for all

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

not b

OpenStudy (anonymous):

not d

OpenStudy (anonymous):

a and c

OpenStudy (er.mohd.amir):

whole numbers starts from 0,1,2,3......................

OpenStudy (anonymous):

b and c

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (anonymous):

c and d

OpenStudy (er.mohd.amir):

why c and d

OpenStudy (anonymous):

and a

OpenStudy (er.mohd.amir):

0-1=-1 which is not in C

OpenStudy (anonymous):

best choice

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d and a

OpenStudy (er.mohd.amir):

among all B is correct

OpenStudy (anonymous):

and d

OpenStudy (anonymous):

not d

OpenStudy (er.mohd.amir):

closed means all output is in given set

OpenStudy (er.mohd.amir):

0-2 = -2 not in set D

OpenStudy (anonymous):

4 some reasons so only b

OpenStudy (er.mohd.amir):

yes only B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. is a negative integer

OpenStudy (anonymous):

d p out of Q

OpenStudy (anonymous):

a and d

OpenStudy (anonymous):

am i correct

OpenStudy (er.mohd.amir):

addition of two -ve numbers is also -ve

OpenStudy (anonymous):

so

OpenStudy (er.mohd.amir):

multipication of two -ve is +ve number

OpenStudy (anonymous):

not b

OpenStudy (er.mohd.amir):

subtraction of two -ve int. may be -ve or +ve

OpenStudy (anonymous):

that is a maybe so no c

OpenStudy (er.mohd.amir):

so A and B are correct

OpenStudy (er.mohd.amir):

complete the D part

OpenStudy (anonymous):

not b is wrong it will be positive

OpenStudy (anonymous):

yes i get it

OpenStudy (anonymous):

Is the set of multiples of 4 closed under addition? Explain why or provide a counterexample if not. A. Yes, because the sum of any two multiples of 4 is also a multiple of 4. B. Yes, because the sum of any two multiples of 4 is also a multiple of 8. C. No, and a counterexample is –4 + (–8) = –12. D. No, and a counterexample is 26 + 16 = 42.

OpenStudy (er.mohd.amir):

if p and q are -ve then p+q is -ve , p*q is +ve , p-q may be +ve or -ve

OpenStudy (er.mohd.amir):

A

OpenStudy (anonymous):

not d

OpenStudy (anonymous):

ok

OpenStudy (er.mohd.amir):

26 is not multiple of 4

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

thanks passed

OpenStudy (anonymous):

i will post a new Q

OpenStudy (er.mohd.amir):

welcome